# Energy dependence of cross sections in proton–proton and antiproton–proton collisions

### v1

@@ IV Fit results and discussion
12 $1.94\pm 0.13$ $8.50\pm 0.08$ $5946.5\pm 1.8$ $39.0\pm 1.9$ $2000\pm 700$ 1.01
0.0 (fixed) $8.4\pm 0.4$ $5926.5\pm 2.3$ $50\pm 4$
13 $2.6\pm 0.4$ $8.5\pm 0.4$ $5945.2\pm 2.3$ $43\pm 14$ $2500\pm 900$ 1.05
0.0 (fixed) $8.6\pm 0.9$ $5930\pm 10$ $42\pm 16$
14 0.0 (fixed) $8.5\pm 0.5$ 6.0 (fixed) $12\pm 6$ $1700\pm 600$ 0.82
–//– $8.5\pm 0.5$ –//– $15\pm 6$
15 0.0 (fixed) $9.7\pm 0.2$ 6.0 (fixed) 0.0 (fixed) 0.0 (fixed) 0.10
–//– $9.44\pm 0.05$ –//– –//–
@@
As it is seen suggested approximation functions (4), (5) provide reasonable fit qualities at smallest $\sqrt{s_{\scriptsize{\mbox{min}},1}}=3$ GeV and, may be, at $\sqrt{s_{\scriptsize{\mbox{min}},2}}=5$ GeV; statistically acceptable $\chi^{2}/\mbox{n.d.f.}$ are obtained at any other $\sqrt{s_{\scriptsize{\mbox{min}}}}$. Thus, as previously Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 the parameterizations for the terms of the set $\mathcal{G}_{1}$ deduced within AQFT show good agreement with available experimental data. The recent RHIC PLB-808-135663-2020 and LHC EPJC-83-441-2023 ; EPJC-79-785-2019 results shown by open symbols in Fig. 1 agree well with general trends for global $pp$ scattering parameters (Fig. 1a, c) especially for $\sigma_{\scriptsize{\mbox{t}}}^{+}$ (Fig. 1a). Therefore, one can expect the new data with respect to the DB24 will negligibly effect on the values of fit parameters, especially for high $s_{\scriptsize{\mbox{min}}}$. This qualitative hypothesis is quite confirmed by many results of our previous studies Okorokov-CPC-46-083105-2022 ; Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 ; Okorokov-PAN-81-508-2018 made for earlier different versions of experimental database for terms of the set $\mathcal{G}_{1}$.
@@
The values of all terms of the set $\mathcal{P}$ from Table 4 obtained for DB24 at low and intermediate $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{7}$ mostly coincide with the values of corresponding fit parameters from Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 within errors for any $s_{\scriptsize{\mbox{min}}}$ indicated above. Fit qualities are (very) close to the published ones Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 at equal $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{7}$. As consequence all features of the fit curves in Fig. 1 shown by solid ($\sqrt{s_{\scriptsize{\mbox{min}},2}}=5$ GeV) and dashed ($\sqrt{s_{\scriptsize{\mbox{min}},7}}=30$ GeV) lines agree with detailed discussion for our previous works Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 . Detailed analysis shows that values of some free parameters coincide with null or / and its asymptotic levels for modified Froissart–Martin limit within uncertainties starting from $\sqrt{s_{\scriptsize{\mbox{min}},10}}=60$ GeV already. In these cases corresponding experimental samples from DB24 are fitted at fixed values of such parameters (Table 2). Thus the functional behavior of $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)$ approaches consecutively and gradually to the modified Froissart–Martin limit with growth of $\sqrt{s_{\scriptsize{\mbox{min}}}}$ starting with high value $\sqrt{s_{\scriptsize{\mbox{min}},14}}=1$ TeV. As it is seen from Table 4 the following relations are valid $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)=\beta_{\pm}\sigma_{\scriptsize{\mbox{FM}}}$ and $\rho^{\pm}(s)=\rho_{\scriptsize{\mbox{FM}}}$ at highest for simultaneous fit $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV. Moreover $\sigma_{\scriptsize{\mbox{t}}}^{+}(s)=\sigma_{\scriptsize{\mbox{t}}}^{-}(s)$ due to coincidence of $\beta_{+}$ and $\beta_{-}$ within 1.04 standard deviation (Table 4). It means that any formulations of Pomeranchuk theorem are valid and all global scattering parameters considered here reach the modified Froissart–Martin limit at least in functional sense because of $\beta_{\pm}\ll\beta_{\scriptsize{\mbox{FM}}}$. The result $\sqrt{s_{a,\scriptsize{\mbox{FM}}}}=5$ TeV can be considered as qualitative indication only for the onset of asymptotic region for global scattering parameters with point of view of the theorems under consideration because of the reasons discussed above, namely, (i) limited data sample for $pp$ collisions, especially, for $\rho^{+}(s)$, (ii) availability of experimental results for $\rho^{+}(s)$ in (very) narrow energy range $\sqrt{s}=7-8$ TeV, and (iii) absence of $\bar{p}p$ data at all. All of these can influence on the robustness and stability for values of terms of the set $\mathcal{P}$ and, consequently, on physical conclusions. On the other hand, the result of the simultaneous fit at $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV agree qualitatively with the previous individual analysis of $\sigma_{\footnotesize\mbox{t}}^{+}(s)$ at $\sqrt{s_{\scriptsize{\mbox{min}},16}}=10$ TeV Okorokov-PAN-81-508-2018 provided the functional behavior of this quality close to the (modified) Froissart–Martin limit.

### v2

@@ IV Fit results and discussion
13 $2.4\pm 0.5$ $8.60\pm 0.12$ $5944.4\pm 2.1$ $40\pm 3$ $2300\pm 700$ 1.17
0.0 (fixed) $8.3\pm 0.4$ $5934\pm 4$ $46\pm 4$
14 0.0 (fixed) $7.1\pm 0.7$ 6000 (fixed) $26\pm 7$ $1800\pm 800$ 1.25
–//– $7.6\pm 1.2$ –//– $22\pm 8$
15 0.0 (fixed) $9.69\pm 0.06$ 6000 (fixed) 0.0 (fixed) 0.0 (fixed) 0.99
–//– $10.3\pm 0.4$ –//– –//–
@@
As it is seen suggested approximation functions (4), (5) provide reasonable fit qualities at smallest $\sqrt{s_{\scriptsize{\mbox{min}},1}}=3$ GeV and, may be, at $\sqrt{s_{\scriptsize{\mbox{min}},14}}=1$ TeV; statistically acceptable $\chi^{2}/\mbox{n.d.f.}$ are obtained at any other $\sqrt{s_{\scriptsize{\mbox{min}}}}$. Thus, as previously Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 the parameterizations for the terms of the set $\mathcal{G}_{1}$ deduced within AQFT show good agreement with available experimental data. The recent RHIC PLB-808-135663-2020 and LHC EPJC-83-441-2023 ; EPJC-79-785-2019 results shown by open symbols in Fig. 1 agree well with general trends for global $pp$ scattering parameters (Fig. 1a, c) especially for $\sigma_{\scriptsize{\mbox{t}}}^{+}$ (Fig. 1a). Therefore, one can expect the new data with respect to the DB24 will negligibly effect on the values of fit parameters, especially for high $s_{\scriptsize{\mbox{min}}}$. This qualitative hypothesis is quite confirmed by many results of our previous studies Okorokov-CPC-46-083105-2022 ; Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 ; Okorokov-PAN-81-508-2018 made for earlier different versions of experimental database for terms of the set $\mathcal{G}_{1}$.
@@
The values of all terms of the set $\mathcal{P}$ from Table 4 obtained for DB24 at low and intermediate $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{7}$ mostly coincide with the values of corresponding fit parameters from Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 within errors for any $s_{\scriptsize{\mbox{min}}}$ indicated above. Fit qualities are (very) close to the published ones Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 at equal $\bigl\{\sqrt{s_{\scriptsize{\mbox{min}},i}}\bigr\}_{i=1}^{7}$. As consequence all features of the fit curves in Fig. 1 shown by solid ($\sqrt{s_{\scriptsize{\mbox{min}},2}}=5$ GeV) and dashed ($\sqrt{s_{\scriptsize{\mbox{min}},7}}=30$ GeV) lines agree with detailed discussion for our previous works Okorokov-IJMPA-25-5333-2010 ; Okorokov-IJMPA-32-1750175-2017 . Detailed analysis shows that values of some free parameters coincide with null or / and its asymptotic levels for modified Froissart–Martin limit within uncertainties starting from $\sqrt{s_{\scriptsize{\mbox{min}},10}}=60$ GeV already. In these cases corresponding experimental samples from DB24 are fitted at fixed values of such parameters (Table 2). Thus the functional behavior of $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)$ approaches consecutively and gradually to the modified Froissart–Martin limit with growth of $\sqrt{s_{\scriptsize{\mbox{min}}}}$ starting with high value $\sqrt{s_{\scriptsize{\mbox{min}},14}}=1$ TeV. As it is seen from Table 4 the following relations are valid $\sigma_{\scriptsize{\mbox{t}}}^{\pm}(s)=\beta_{\pm}\sigma_{\scriptsize{\mbox{FM}}}$ and $\rho^{\pm}(s)=\rho_{\scriptsize{\mbox{FM}}}$ at highest boundary for simultaneous fit $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV. Moreover $\sigma_{\scriptsize{\mbox{t}}}^{+}(s)=\sigma_{\scriptsize{\mbox{t}}}^{-}(s)$ due to coincidence of $\beta_{+}$ and $\beta_{-}$ within better than 1.33 standard deviation (Table 4). It means that any formulations of Pomeranchuk theorem are valid and all global scattering parameters considered here reach the modified Froissart–Martin limit at least in functional sense because of $\beta_{\pm}\ll\beta_{\scriptsize{\mbox{FM}}}$. The result $\sqrt{s_{a,\scriptsize{\mbox{FM}}}}=5$ TeV can be considered as qualitative indication only for the onset of asymptotic region for global scattering parameters with point of view of the theorems under consideration because of the reasons discussed above, namely, (i) limited data sample for $pp$ collisions, especially, for $\rho^{+}(s)$, (ii) availability of experimental results for $\rho^{+}(s)$ in (very) narrow energy range $\sqrt{s}=7-8$ TeV, and (iii) absence of $\bar{p}p$ data at all. All of these can influence on the robustness and stability for values of terms of the set $\mathcal{P}$ and, consequently, on physical conclusions. On the other hand, the result of the simultaneous fit at $\sqrt{s_{\scriptsize{\mbox{min}},15}}=5$ TeV agree qualitatively with the previous individual analysis of $\sigma_{\footnotesize\mbox{t}}^{+}(s)$ at $\sqrt{s_{\scriptsize{\mbox{min}},16}}=10$ TeV Okorokov-PAN-81-508-2018 provided the functional behavior of this quality close to the (modified) Froissart–Martin limit.